# Partial-immunity of two-photon correlation against wavefront distortion for spatially entangled photons

### v1

@@ Acknowledgements.
We acknowledge the TIFR HPC facility used for the numerical simulations. We express our gratitude to the Department of Atomic Energy, Government of India, for funding for Project Identification No. RTI4002 under DAE OM No. 1303/1/2020/R&D-II/DAE/5567, Ministry of Science and Technology, India. The authors declare no conflict of interest.
@@ Appendix A Deformable Mirror (DM) and Zernike polynomials
In this section, we describe the structure of the DM used in the experiment. The DM features a continuous reflective surface with a total mirror diameter of $12.7~{}mm$ and a pupil diameter of $10~{}mm$. The deformation of the mirror surface is controlled by $40$ independent piezoelectric actuators positioned on the back side of the reflective surface, as illustrated in Fig. 7(a). The center of each segment within the circular disc of the DM indicates the position of an actuator. Although these actuators can in principle be controlled independently, any change in the voltage of one actuator is observed to affect its boundary with neighboring segments compromising the independence of the individual pixels.
@@ Appendix B Measurement of Schmidt number
Figure 8: (a) The original beam recorded using EMCCD camera in the far-field. (b) Two-photon momentum anti-correlation.
@@
$\rho(k_{1})\propto e^{-k_{1}^{2}\frac{4\sigma_{+}^{2}\sigma_{-}^{2}}{\sigma_{+}^{2}-\sigma_{-}^{2}}}$ (11)
@@
To experimentally obtain $\sigma_{m}$ and $\sigma_{k}$, the measurements need to be performed in the far-field (momentum space), which can be obtained by inserting a lens after the nonlinear crystal and detecting the photons in the back focal plane. Fig. 8 (a) shows the beam captured using an EMCCD which gives the marginal distribution of a photon and Fig. 8 (b) shows the projection of JPD in $k_{1}+k_{2}$ space obtained by taking $3000$ frames using EMCCD and post-analyzing using the method discussed in Reference Chatterjee et al. (2024). The values of $\sigma_{m}$ and $\sigma_{j}$ obtained by 2D Gaussian fitting are $121.97$ and $4.59$ pixels respectively. Therefore the Schmidt number of SPDC photons is $707\pm 10$.
@@ Appendix C Comparing the interference pattern for two wavelengths
Here, we compare the speckle patterns corresponding to two wavelengths, $405~{}nm$ and $810~{}nm$, for identical configurations of the DM. A collimated coherent light source with a waist size comparable to the DM’s radius is used to illuminate the DM at both wavelengths, and the interference patterns are observed in the far field. The results are shown in Fig. 9. Due to the wavelength dependence of the phase for a given path length, the $405~{}nm$ light experiences double the phase shift compared to the $810~{}nm$ light. Consequently, the speckle patterns for the two wavelengths are not expected match. The strength of disorder introduced by the DM can be quantified using the speckle contrast, defined as:
@@
Figure 9: Interference pattern observed with coherent light source of wavelength (a) $405~{}nm$ and (b) $810~{}nm$ for same configuration of the DM.
where $\sigma_{I}$ represents the standard deviation of the intensity distribution, and $\langle I\rangle$ denotes the mean intensity. This measure provides a quantitative comparison of the disorder for the two wavelengths. The speckle contrast values for the two wavelengths ($405~{}nm$ and $810~{}nm$) are $0.89$ and $0.96$ respectively, indicating that the disorder strength is comparable for both.

### v2

@@ Acknowledgements.
We acknowledge the TIFR CC HPC facility used for the numerical simulations. We express our gratitude to the Department of Atomic Energy, Government of India, for funding for Project Identification No. RTI4002 under DAE OM No. 1303/1/2020/R&D-II/DAE/5567, Ministry of Science and Technology, India. The authors declare no conflict of interest.
@@ Appendix A Deformable Mirror (DM) and Zernike polynomials
In this section, we describe the structure of the DM used in the experiment. The DM features a continuous reflective surface with a total mirror diameter of $12.7$ mm and a pupil diameter of $10$ mm. The deformation of the mirror surface is controlled by $40$ independent piezoelectric actuators positioned on the back side of the reflective surface, as illustrated in Fig. 7(a). The center of each segment within the circular disc of the DM indicates the position of an actuator. Although these actuators can in principle be controlled independently, any change in the voltage of one actuator is observed to affect its boundary with neighboring segments compromising the independence of the individual pixels.
@@ Appendix B Measurement of Schmidt number
Figure 8: (a) The original beam recorded using EMCCD camera in the far field. (b) Two-photon momentum anti-correlation.
@@
$\rho(k_{1})\propto e^{-k_{1}^{2}\frac{4\sigma_{+}^{2}\sigma_{-}^{2}}{\sigma_{+}^{2}+\sigma_{-}^{2}}}$ (11)
@@
To experimentally obtain $\sigma_{m}$ and $\sigma_{j}$, the measurements need to be performed in the far field (momentum space), which can be obtained by inserting a lens after the nonlinear crystal and detecting the photons in the back focal plane. Fig. 8 (a) shows the beam captured using an EMCCD which gives the marginal distribution of a photon and Fig. 8 (b) shows the projection of JPD in $k_{1}+k_{2}$ space obtained by taking $3000$ frames using EMCCD and post-analyzing using the method discussed in Reference Chatterjee et al. (2024). The values of $\sigma_{m}$ and $\sigma_{j}$ obtained by 2D Gaussian fitting are $121.97$ and $4.59$ pixels respectively. Therefore the Schmidt number of SPDC photons is $707\pm 10$.
@@ Appendix C Comparing the interference pattern for two wavelengths
Here, we compare the speckle patterns corresponding to two wavelengths, $405$ nm and $810$ nm, for identical configurations of the DM. A collimated coherent light source with a waist size comparable to the DM’s radius is used to illuminate the DM at both wavelengths, and the interference patterns are observed in the far field. The results are shown in Fig. 9. Due to the wavelength dependence of the phase for a given path length, the $405$ nm light experiences double the phase shift compared to the $810$ nm light. Consequently, the speckle patterns for the two wavelengths are not expected match. The strength of disorder introduced by the DM can be quantified using the speckle contrast, defined as:
@@
Figure 9: Interference pattern observed with coherent light source of wavelength (a) $405$ nm and (b) $810$ nm for same configuration of the DM.
where $\sigma_{I}$ represents the standard deviation of the intensity distribution, and $\langle I\rangle$ denotes the mean intensity. This measure provides a quantitative comparison of the disorder for the two wavelengths. The speckle contrast values for the two wavelengths ($405$ nm and $810$ nm) are $0.89$ and $0.96$ respectively, indicating that the disorder strength is comparable for both.